Explain button cells and also give the reactions at the anode and the cathode .
1
Expert's answer
2019-11-18T06:28:44-0500

Solution.

The button cell is a small separate battery cell in the shape of a squat cylinder. The metal forms the lower part of the body and the positive end of the cell (cathode). The insulated top cover is the negative terminal (anode). MnO2, Ag2O, CuO can act as a cathode. As the anode can act Zn, Li. For example, MnO2 - Zn, CuO - Li.

MnO2 (cathode) - Zn(anode)

Cathode reaction:

"2MnO2 + 2H3O^+ + 2e = 2MnO(OH) + 2H2O"

Anode reaction:

"Zn^0 - 2e = Zn^{2+}"

CuO(cathode) - Li(anode)

Cathode reaction:

"CuO + H2O+2e = Cu^0 + 2OH^-"

Anode reaction:

"Li^0 - e = Li^+"

Answer:

The button cell is a small separate battery cell in the shape of a squat cylinder. The metal forms the lower part of the body and the positive end of the cell (cathode). The insulated top cover is the negative terminal (anode). MnO2, Ag2O, CuO can act as a cathode. As the anode can act Zn, Li. For example, MnO2 - Zn, CuO - Li.

MnO2 (cathode) - Zn(anode)

Cathode reaction:

"2MnO2 + 2H3O^+ + 2e = 2MnO(OH) + 2H2O"

Anode reaction:

"Zn^0 - 2e = Zn^{2+}"

CuO(cathode) - Li(anode)

Cathode reaction:

"CuO + H2O+2e = Cu^0 + 2OH^-"

Anode reaction:

"Li^0 - e = Li^+"